OpenStudy (anonymous):

simplify 4rad3+7rad3-2rad3

OpenStudy (anonymous):

9rad3

OpenStudy (anonymous):

how about 2rad3+rad12

OpenStudy (anonymous):

4rad3

OpenStudy (anonymous):

throw in as many as u want

OpenStudy (anonymous):

3rad8+7rad2

OpenStudy (anonymous):

13rad2

OpenStudy (anonymous):

9rad4-12rad4

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

+6 what

OpenStudy (anonymous):

-6

OpenStudy (anonymous):

rad12

OpenStudy (anonymous):

2rad3

OpenStudy (anonymous):

ar u done?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\[4 \sqrt{3}+7 \sqrt{3}-2 \sqrt{3}=9 \sqrt{3} \]Factor out the rad 3 and add up the integers. @Ishaan94 was the first to post the result.

OpenStudy (anonymous):

rad5+4rad5

OpenStudy (anonymous):

5rad5

OpenStudy (anonymous):

u can do em on ur own

OpenStudy (anonymous):

kinda

OpenStudy (anonymous):

so y r u asking these?

OpenStudy (anonymous):

cuz i did the work im just checking if there right

OpenStudy (anonymous):

u don't need such method to check....

OpenStudy (anonymous):

u gotta be confident on ur answers ... u aint getting solutions for every problem

OpenStudy (anonymous):

ur right

OpenStudy (anonymous):

so r u going to stop these things?

OpenStudy (anonymous):

i will

OpenStudy (anonymous):

ask only the ones which r difficult... and try understand method.. i'l help you I promise

OpenStudy (anonymous):

so tell me the most difficult ones

OpenStudy (anonymous):

ok this problem i suck at 8rad3-rad48 im lost , i dont like math

OpenStudy (anonymous):

would it be 8 three times

OpenStudy (anonymous):

Its Okay, Tons of ppl don't like Maths Now To The Problem 8rad3 - rad 6x8 8rad3 - 4rad 3 ; 6= 2x3 ; 8 = 2x2x2 rad3(8-4) 4rad3

OpenStudy (anonymous):

wow

OpenStudy (anonymous):

yeah wow

OpenStudy (anonymous):

6*8 equal 48

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

how did you get 4rad 3

OpenStudy (anonymous):

rad ( 48) rad(2x2x2x2x3) -------- ^ | see these 2x2x2x2 these come out as 2x2 =4 so in problem rad3(8-4) 4rad3

OpenStudy (anonymous):

oo okay gotcha

OpenStudy (anonymous):

Nice :)

OpenStudy (anonymous):

thank you so much for taking the time :)

OpenStudy (anonymous):

Pleasure's Myne